# Double-quantum-dot Andreev molecules: Phase diagrams and critical evaluation of effective models

### v1

@@ Abstract
Electrostatically defined quantum dots (QDs) in nanowires or carbon nanotubes, coupled with superconducting (SC) environments, offer a promising pathway for constructing quantum electronic circuits with diverse functionalities [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19]. These systems are characterized by the emergence of states within the SC gap which are referred to as either Yu-Shiba-Rusinov (YSR) or Andreev bound states (ABS) [20, 21, 22], which are influenced by strong Coulomb interactions arising from their underlying electronic structure. Consequently, a rich phenomenology of quantum phases is observed [23, 24].
@@
While early efforts focused on single QD devices, advancements in fabrication have enabled the development of double-quantum-dot (DQD) systems in both serial and parallel configurations. Initially, parallel configurations predominantly appeared as central components of Cooper pair splitting (CPS) devices [25, 26, 27, 28, 29, 30, 31, 32, 33], correspondingly driving the bulk of the theoretical research into this particular direction [34, 35, 36, 37, 38, 39, 40, 41, 42]. However, coupling two sites to a common reservoir following the Bardeen–Cooper–Schrieffer (BCS) theory has broader and more fundamental significance, as highlighted in recent proposal of quartet superconductivity at equilibrium [43] or in qubit engineering [44, 45]. A central concept is then the formation of Andreev molecular states, which have a non-local spatial nature, as recently studied experimentally in [46, 47].
@@
Due to the strong electronic interactions various competing effects emerge in DQD devices at once. In SC serial configurations, Kondo screening and superconductivity shape simultaneously the outcome. Parallel DQD configurations have on top an inherent potential of an additional third effect, namely the lead-mediated interactions between the constituent QDs [48]. In this regard, two opposing scenarios are particularly notable. If the spatial distance between the two QDs is negligible compared to the SC coherence length $\zeta$ of their common lead, one may assume that a single screening channel is available to both QDs. Thus, a theoretical model of one BCS lead connected simultaneously to two QDs may be formulated, as schematically shown in Fig. 1(a). In the opposite limit, one may imagine the two QDs being physically far apart. Then, even if connected to a common lead, they will practically experience two independent screening channels already in a metallic case [48]. To model such a situation with superconductivity present, two BCS channels, each coupled to a separate QD, as shown in Fig. 1(b), are required. Consequently, all lead-mediated interactions between QDs disappear and a relatively well-understood serial DQD scenario emerges effectively. Here, an extensive theoretical understanding already exists, which encompasses accurate non-perturbative results obtained via the Numerical Renormalization Group (NRG) or Quantum Monte Carlo (QMC) methods [49, 50, 51] [^1] If the study is limited only to the subgap properties, the recently developed surrogate model [52] might also be applied.. The accumulated systematic knowledge even allowed the development of numerically inexpensive effective descriptions via simplistic models, such as zero bandwidth (ZBW) [53, 54], atomic limit (AL) [55, 56, 57] and its generalized version (GAL) [51].
@@
In the one-SC-channel scenario, the current understanding is considerably less systematic [44, 58], with effective models often substituting the full descriptions, despite their extensions being not sufficiently tested. Using NRG [59, 60, 61, 62] as our primary method, we therefore systematically solve the parallel DQD case across a wide parameter space, focusing particularly on regions with significant lead-mediated interactions between the QDs. The phase diagrams we find include singlet, doublet, and relatively uncommon triplet ground states (GSs), with the latter one being an important signature of the presence of significant lead-mediated interactions. Finally, we show that triplet phases cannot be reproduced by neither AL nor ZBW models, revealing thus their critical failure even in terms of qualitative predictions. This consequently underscores the necessity for increased caution and thorough cross-verification with either exact or well-defined approximate methods to avoid misleading conclusions when extrapolating AL or ZBW models to more complex QD systems.
@@ I.1 SC-Anderson model for the parallel DQD scenario
Our objective is to model a DQD screened by one spinfull SC channel, as shown schematically in Fig. 1$(a)$. We therefore consider an SC Anderson model with two QDs coupled to the same reservoir. The Hamiltonians for the upper ($j=1$) and lower QDs ($j=2)$ are then given as
@@
with $t$ being the corresponding inter-dot hopping matrix elements. Without sake of generality, we assume that the inter-dot charging energy is negligible here. Moreover, we also assume $U_{1}=U_{2}\equiv U$ and $\Gamma_{1}=\Gamma_{2}\equiv\Gamma$. Finally, the SC lead is described via the ordinary BCS theory as
@@
with $c^{\vphantom{\dagger}}_{\mathbf{k}\sigma}$ ($c^{\dagger}_{\mathbf{k}\sigma}$) being the annihilation (creation) operator of the electron with spin $\sigma$ and momentum $\mathbf{k}$ in the SC lead. Using the BCS gauge invariance, zero phase is selected in the lead.
@@ I.2 The tunneling self-energy
$\mathbb{G}_{0}(\omega^{+})$ $=$ $\left[\omega^{+}\mathbb{1}-\left(\begin{array}[]{c c}\mathbb{E}_{11}&\mathbb{E}_{12}\\ \mathbb{E}_{21}&\mathbb{E}_{22}\end{array}\right)-\left(\begin{array}[]{c c}\mathbb{\Sigma}_{11}(\omega^{+})&\mathbb{\Sigma}_{12}(\omega^{+})\\ \mathbb{\Sigma}_{21}(\omega^{+})&\mathbb{\Sigma}_{22}(\omega^{+})\end{array}\right)\right]^{-1},$ (19)
@@
[…]
@@ I.4 Zero bandwidth approximations
We therefore rigorously explore the potential of ZBW approximation to model the present parallel DQD setup given by the full Hamiltonian (5). To this end, we first lay down its ZBW approximation. Following Ref. [54], we drop the $\mathbf{k}$-dependence in the BCS and tunneling Hamiltonians, leaving $H_{d,j}$ and $H_{t}$ unaltered. We label the orbital representing the lead as 00 and define its corresponding electron annihilation (creation) operator as $c^{\vphantom{\dagger}}_{0\sigma}$ ($c^{\dagger}_{0\sigma}$). The total Hamiltonian in the ZBW approximation becomes then
@@
Figure 2: The NRG results for the phase diagrams for the DQD system of two identical QDs ($U_{1}=U_{2}\equiv U$), as functions of the detuning parameters $-\delta_{1}/U$ and $-\delta_{2}/U$ for different values of the superconducting energy gap $\Delta/U$, as indicated. Coupling-symmetric scenario $\Gamma_{1}/U=\Gamma_{2}/U=\equiv\Gamma/U=0.05$ maximizing lead-mediated correlations ($\nu=1$) is selected. The white, yellow, and green regions correspond to singlet ($S=0$), doublet ($S=1/2$), and triplet ($S=1$) GS, respectively. (a) In the absence of superconductivity ($\Delta=0$), non-zero cross-correlations mediated by the common lead result in a doublet GS island at the center of the phase diagram. (b)-(c) As $\Delta/U$ increases from $5\times 10^{-5}$ to $10^{-4}$, the doublet GS island elongates into a cross-shaped region. (d)-(e) For $\Delta/U=5\times 10^{-4}$, a triplet GS emerges in the center of the phase diagram. (f)-(g) The cross-shaped doublet region splits into two separate stripes along the main diagonal $\delta_{1}=\delta_{2}$, and the triplet region begins to shrink. (h) As $\Delta/U$ increases further to $0.5$, the triplet GS disappears, and a striped pattern of singlet and doublet phases appears in the direction of the main diagonal $\delta_{1}=\delta_{2}$. All calculations have been performed via NRG with discretization parameter $\Lambda=2$ exploiting full spin symmetry with at least $1000$ multiplets kept during calculations, assuming $U=0.2D$ and the band halfwidth $D\equiv 1$ used as energy unit.
@@ II.1 Results in the absence of inter-dot hopping
Figure 3: The NRG results for the phase diagrams for the coupling symmetric system of two identical parallel QDs at $\nu=1$ as a function of the detuning parameters $-\delta_{1}/U$ and $-\delta_{2}/U$ for varying superconducting gap $\Delta/U$. The white, yellow, and green regions correspond to the singlet ($S=0$), doublet ($S=1/2$), and triplet ($S=1$) GS, respectively. (a1)-(a5) For $\Gamma/U=0.1$, the evolution with increasing $\Delta/U$ is qualitatively similar to Fig. 2, where a cross-shaped doublet region splits into separate stripes, and a triplet GS emerges and then disappears as $\Delta/U$ increases. (b1)-(b5) For $\Gamma/U=0.2$, the Kondo screening effects dominate over the lead-mediated correlations, preventing the formation of a triplet GS for any value of $\Delta/U$. (c1)-(c5) At a stronger coupling of $\Gamma/U=0.5$, the Kondo screening processes dominate the behavior, resulting in phase diagrams with only centrally-positioned doublet region immersed in a singlet GS area. The other parameters are the same as in Fig. 2.
@@
We first examine the $\Gamma/U=0.05$ case presented in Fig. 2. When the SC effects are excluded, as in Fig. 2$(a)$ for $\Delta=0$, the phase diagram is governed solely by the Kondo screening and RKKY interactions [48], leading to a well-known behavior of a doublet phase in the center of the diagram occupying almost the entire region of $-1/2<\delta_{1},\delta_{2}<1/2$, in line with the qualitative picture of Fig. 1(c). The small bendings of the observed parity transition lines are to be attributed to the presence of lead-mediated RKKY interactions for finite coupling $\Gamma$. The doublet GS in the middle of the phase diagram is associated with the underscreened Kondo effect [74, 75, 76, 77], in which only one of the spins occupying the DQD can be screened, while the other one is responsible for $S=1/2$ GS.
@@
Let us finally note that so far all observed phase diagrams possessed two symmetric mirror axes $\delta_{1}=\delta_{2}$ and $\delta_{1}=-\delta_{2}$, respectively. While the first is a trivial consequence of selecting coupling symmetric scenario with two identical dots, the latter one is the consequence of the ph-transformations at $t=0$, as discussed in Sec. V.1. However, when $t\neq 0$, this second symmetry will be broken as shown below.
@@ II.2 Effects of non-zero inter-dot hopping
The second most important feature is the gradual disruption of the triplet GS as $t/U$ increases. This evolution is expected due to inter-dot hopping mediating fluctuations between the two constituent dots, favoring singlet over triplet GSs. The triplet area starts to shrink and move apart in the quadrant defined by $-\delta_{1},-\delta_{2}>0$ when $t>0$, which is accompanied by the contraction of the doublet area, favoring the singlet GS once again. Eventually, the doublet regions separate then completely as seen in Figs. 4(b3), (c3), and (d3). Finally, at $t/U=0.25$ the doublet GS are completely removed from the depicted portion of the $-\delta_{1},-\delta_{2}>0$ quadrant with only the case of $\Delta/U=0.05$ being a notable exception. Contrary, in the quadrant $-\delta_{1},-\delta_{2}<0$, a connected stripe of doublet GS remains intact in the area of $-\delta_{1},-\delta_{2}<0$.
@@ III.1 Limitations of the AL theory
Figure 5: Phase diagrams of two identical parallel dots ($U_{1}=U_{2}\equiv U$) calculated using the AL theory for the symmetric coupling scenario with $\Gamma^{\mathrm{AL}}_{1}=\Gamma^{\mathrm{AL}}_{2}\equiv\Gamma^{\mathrm{AL}}$ at $\nu=1$ as a function of the detuning parameters, $-\delta_{1}/U$ and $-\delta_{2}/U$. At $t=0$, $\Gamma^{\mathrm{AL}}/U$ remains the only relevant scale to be varied in the AL theory. Panels (a)-(c) with $\Gamma^{\mathrm{AL}}/U\leq 0.25$ are in a good qualitative agreement with the NRG results observed for large $\Delta$ in Figs. 2 and 3. However, the remaining panels (d)-(f), with larger $\Gamma^{\mathrm{AL}}/U$ values, have no counterpart in the exact results. Overall, the AL fails in predicting the triplet GS completely. The same also applies to GAL, as it only rescales the parameters that enter the AL theory.

### v2

@@ Abstract
Electrostatically defined quantum dots (QDs) in nanowires or carbon nanotubes, coupled with superconducting (SC) environments, offer a promising pathway for constructing quantum electronic circuits with diverse functionalities [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19]. These systems are characterized by the emergence of states within the SC gap which are referred to as either Yu-Shiba-Rusinov (YSR) or Andreev bound states (ABS) [20, 21, 22], which are influenced by strong Coulomb interactions arising from their underlying electronic structure. Consequently, a rich phenomenology of quantum phases is observed [23, 24, 25].
@@
While early efforts focused on single QD devices, advancements in fabrication have enabled the development of double-quantum-dot (DQD) systems in both serial and parallel configurations. Initially, parallel configurations predominantly appeared as central components of Cooper pair splitting (CPS) devices [26, 27, 28, 29, 30, 31, 32, 33, 34], correspondingly driving the bulk of the theoretical research into this particular direction [35, 36, 37, 38, 39, 40, 41, 42, 43]. However, coupling two sites to a common reservoir following the Bardeen–Cooper–Schrieffer (BCS) theory has broader and more fundamental significance, as highlighted in recent proposal of quartet superconductivity at equilibrium [44], in qubit engineering [45, 46] or in minimal Kitaev chains [47, 48]. A central concept is then the formation of Andreev molecular states, which have a non-local spatial nature, as recently studied experimentally in [49, 50, 51, 52].
Due to the strong electronic interactions various competing effects emerge in DQD devices at once. In SC serial configurations, Kondo screening and superconductivity shape simultaneously the outcome. Parallel DQD configurations have on top an inherent potential of an additional third effect, namely the lead-mediated interactions between the constituent QDs [53]. In this regard, two opposing scenarios are particularly notable.
If the spatial distance between the two QDs is negligible compared to the SC coherence length $\zeta$ of their common lead, one may assume that a single screening channel is available to both QDs. Thus, a theoretical model of one BCS lead connected simultaneously to two QDs may be formulated, as schematically shown in Fig. 1(a). In the opposite limit, one may imagine the two QDs being physically far apart. Then, even if connected to a common lead, they will practically experience two independent screening channels already in the metallic case [53]. Consequently, two BCS channels are required in the presence of superconducting correlations in the leads. They are each coupled to a separate QD, as shown in Fig. 1(b).
@@
In the latter scenario, all lead-mediated interactions between QDs disappear and a relatively well-understood serial DQD scenario emerges effectively. Here, an extensive theoretical understanding already exists, which encompasses accurate non-perturbative results obtained via the Functional and Numerical Renormalization Group (NRG) or Quantum Monte Carlo (QMC) methods [54, 55, 56, 57] [^1] If the study is limited only to the subgap properties, the recently developed surrogate model [58] might also be applied.. The accumulated systematic knowledge even allowed the development of numerically inexpensive effective descriptions via simplistic models, such as zero bandwidth (ZBW) [59, 60], atomic limit (AL) [61, 62, 63] and its generalized version (GAL) [56]. While in ZBW, the $\mathbf{k}$-dependency of the lead is dropped, describing effectively all screening effects via single orbital, AL can be viewed as a $\Delta\rightarrow\infty$ description. Adding proper parameter rescallings in AL leads then to the GAL theory.
@@
In the one-SC-channel scenario, the current understanding is considerably less systematic [45, 64, 65], with effective models often substituting the full descriptions, despite their extensions not being sufficiently tested. Using NRG [66, 67, 68, 69, 70, 71] as our primary method, we therefore systematically solve the parallel DQD case across a wide parameter space, focusing particularly on regions with significant lead-mediated interactions between the QDs.
The phase diagrams we find include singlet, doublet, and relatively uncommon triplet ground states (GSs) [57], with the latter one being an important signature of the presence of significant lead-mediated interactions. Finally, we show that triplet phases can be reproduced by neither AL nor ZBW models, revealing thus their critical failure even in terms of qualitative predictions. This consequently underscores the necessity for increased caution and thorough cross-verification with either exact or well-defined approximate methods to avoid misleading conclusions when extrapolating AL or ZBW models to more complex QD systems.
@@ I.1 SC-Anderson model for the parallel DQD scenario
Our objective is to model a DQD screened by one spinfull SC channel, as shown schematically in Fig. 1(a)(a). We therefore consider an SC Anderson model with two QDs coupled to the same reservoir. The Hamiltonians for the upper ($j=1$) and lower QDs ($j=2)$ are then given as
@@
with $t$ being the corresponding inter-dot hopping matrix elements. Here, we assume that the inter-dot charging energy is negligible. Moreover, we also assume $U_{1}=U_{2}\equiv U$ and $\Gamma_{1}=\Gamma_{2}\equiv\Gamma$. Finally, the SC lead is described via the ordinary BCS theory as
@@
with $c^{\vphantom{\dagger}}_{\mathbf{k}\sigma}$ ($c^{\dagger}_{\mathbf{k}\sigma}$) being the annihilation (creation) operator of the electron with spin $\sigma$, momentum $\mathbf{k}$, and energy $\varepsilon_{k}$ in the SC lead, while $\Delta$ stands for the gap parameter with a zero phase selected by using the BCS gauge invariance.
@@ I.2 The tunneling self-energy
$\mathbb{G}_{0}(\omega^{+})\hskip-4.2679pt$ $=$ $\hskip-4.2679pt\left[\omega^{+}\mathbb{1}\hskip-1.42262pt-\hskip-1.42262pt\left(\hskip-2.84526pt\begin{array}[]{c c}\mathbb{E}_{11}&\mathbb{E}_{12}\\ \mathbb{E}_{21}&\mathbb{E}_{22}\end{array}\hskip-2.84526pt\right)\hskip-1.42262pt-\hskip-1.42262pt\left(\hskip-2.84526pt\begin{array}[]{c c}\mathbb{\Sigma}_{11}(\omega^{+})&\mathbb{\Sigma}_{12}(\omega^{+})\\ \mathbb{\Sigma}_{21}(\omega^{+})&\mathbb{\Sigma}_{22}(\omega^{+})\end{array}\hskip-2.84526pt\right)\right]^{-1},$ (19)
@@
$\mathbb{\Sigma}_{ij}(\omega^{+})$ $=\frac{\Gamma_{ij}}{\pi}\int_{-D}^{D}\frac{\omega\mathbb{1}-\Delta\sigma_{x}}{\omega^{2}-\Delta^{2}-\varepsilon^{2}+i\eta\,\mathrm{sgn}(\omega)}d\varepsilon$
@@ I.4 Zero bandwidth approximations
We therefore rigorously explore the potential of ZBW approximation to model the present parallel DQD setup given by the full Hamiltonian (5). To this end, we first lay down its ZBW approximation. Following Ref. [60], we drop the $\mathbf{k}$-dependence in the BCS and tunneling Hamiltonians, leaving $H_{d,j}$ and $H_{t}$ unaltered. We label the orbital representing the lead as 0 and define its corresponding electron annihilation (creation) operator as $c^{\vphantom{\dagger}}_{0\sigma}$ ($c^{\dagger}_{0\sigma}$). The total Hamiltonian in the ZBW approximation becomes then
@@
Figure 2: The NRG results for the phase diagrams for the DQD system of two identical QDs ($U_{1}=U_{2}\equiv U$), as functions of the detuning parameters $\delta_{1}/U$ and $\delta_{2}/U$ for different values of the superconducting energy gap $\Delta/U$, as indicated. Coupling-symmetric scenario $\Gamma_{1}/U=\Gamma_{2}/U\equiv\Gamma/U=0.05$ maximizing lead-mediated correlations ($\nu=1$) is selected. The white, yellow, and green regions correspond to singlet ($S=0$), doublet ($S=1/2$), and triplet ($S=1$) GS, respectively. (a) In the absence of superconductivity ($\Delta=0$), non-zero cross-correlations mediated by the common lead result in a doublet GS island at the center of the phase diagram. (b)-(c) As $\Delta/U$ increases from $5\times 10^{-5}$ to $10^{-4}$, the doublet GS island elongates into a cross-shaped region. (d)-(e) For $\Delta/U=5\times 10^{-4}$, a triplet GS emerges in the center of the phase diagram. (f)-(g) The cross-shaped doublet region splits into two separate stripes along the main diagonal $\delta_{1}=\delta_{2}$, and the triplet region begins to shrink. (h) As $\Delta/U$ increases further to $0.5$, the triplet GS disappears, and a striped pattern of singlet and doublet phases appears in the direction of the main diagonal $\delta_{1}=\delta_{2}$. All calculations have been performed via NRG with discretization parameter $\Lambda=2$ exploiting full spin symmetry with at least $1000$ multiplets kept during calculations, assuming $U=0.2D$ and the band halfwidth $D\equiv 1$ used as energy unit.
@@ II.1 Results in the absence of inter-dot hopping
Figure 3: The NRG results for the phase diagrams for the coupling symmetric system of two identical parallel QDs at $\nu=1$ as a function of the detuning parameters $\delta_{1}/U$ and $\delta_{2}/U$ for varying superconducting gap $\Delta/U$. The white, yellow, and green regions correspond to the singlet ($S=0$), doublet ($S=1/2$), and triplet ($S=1$) GS, respectively. (a1)-(a5) For $\Gamma/U=0.1$, the evolution with increasing $\Delta/U$ is qualitatively similar to Fig. 2, where a cross-shaped doublet region splits into separate stripes, and a triplet GS emerges and then disappears as $\Delta/U$ increases. (b1)-(b5) For $\Gamma/U=0.2$, the Kondo screening effects dominate over the lead-mediated correlations, preventing the formation of a triplet GS for any value of $\Delta/U$. (c1)-(c5) At a stronger coupling of $\Gamma/U=0.5$, the Kondo screening processes dominate the behavior, resulting in phase diagrams with only centrally-positioned doublet region immersed in a singlet GS area. The other parameters are the same as in Fig. 2.
@@
We first examine the $\Gamma/U=0.05$ case presented in Fig. 2. When the SC effects are excluded, as in Fig. 2(a)(a) for $\Delta=0$, the phase diagram is governed solely by the Kondo screening and RKKY interactions [53], leading to a well-known behavior of a doublet phase in the center of the diagram occupying almost the entire region of $1/2<\delta_{1},\delta_{2}<1/2$, in line with the qualitative picture of Fig. 1(c). The small bendings of the observed parity transition lines are to be attributed to the presence of lead-mediated RKKY interactions for finite coupling $\Gamma$. The doublet GS in the middle of the phase diagram is associated with the underscreened Kondo effect [82, 83, 84, 85], in which only one of the spins occupying the DQD can be screened, while the other one is responsible for $S=1/2$ GS.
@@
Let us finally note that so far all observed phase diagrams possessed two symmetric mirror axes $\delta_{1}=\delta_{2}$ and $\delta_{1}=-\delta_{2}$, respectively. While the first is a trivial consequence of selecting coupling symmetric scenario with two identical dots, the latter one is the consequence of the ph-transformations at $t=0$, as discussed in Sec. V.1. However, when $t\neq 0$, this second symmetry will be broken as shown next.
@@ II.2 Effects of non-zero inter-dot hopping
The second most important feature is the gradual disruption of the triplet GS as $t/U$ increases. This evolution is expected due to inter-dot hopping mediating fluctuations between the two constituent dots, favoring singlet over triplet GSs. The triplet area starts to shrink and move apart in the quadrant defined by $\delta_{1},-\delta_{2}>0$ when $t>0$, which is accompanied by the contraction of the doublet area, favoring the singlet GS once again. Eventually, the doublet regions separate then completely as seen in Figs. 4(b3), (c3), and (d3). Finally, at $t/U=0.25$ the doublet GS are completely removed from the depicted portion of the $\delta_{1},-\delta_{2}>0$ quadrant with only the case of $\Delta/U=0.05$ being a notable exception. Contrary, in the quadrant $\delta_{1},-\delta_{2}<0$, a connected stripe of doublet GS remains intact in the area of $\delta_{1},-\delta_{2}<0$.
@@ III.1 Limitations of the AL theory
Figure 5: Phase diagrams of two identical parallel dots ($U_{1}=U_{2}\equiv U$) calculated using the AL theory for the symmetric coupling scenario with $\Gamma^{\mathrm{AL}}_{1}=\Gamma^{\mathrm{AL}}_{2}\equiv\Gamma^{\mathrm{AL}}$ at $\nu=1$ as a function of the detuning parameters, $\delta_{1}/U$ and $\delta_{2}/U$. At $t=0$, $\Gamma^{\mathrm{AL}}/U$ remains the only relevant scale to be varied in the AL theory. Panels (a)-(c) with $\Gamma^{\mathrm{AL}}/U\leq 0.25$ are in a good qualitative agreement with the NRG results observed for large $\Delta$ in Figs. 2 and 3. However, the remaining panels (d)-(f), with larger $\Gamma^{\mathrm{AL}}/U$ values, have no counterpart in the exact results. Overall, the AL fails in predicting the triplet GS completely. The same also applies to GAL, as it only rescales the parameters that enter the AL theory.